rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_red
icon_green
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_green
icon_blue
icon_orange
icon_green
icon_blue
icon_blue
icon_red
icon_red
icon_green
icon_red
icon_orange
icon_blue
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_green
 

FS#38536 — Optical line SBG <>ParisGSW

Attached to Project— Reseau Internet et Baies
Maintenance
Backbone EU
CLOSED
100%
We are scheduling a planned work on one of our 2 optical lines Strasbourg <> Paris GSW
- May 21st from 23:00 CET

These operations are intended to change optical parameters on the line in order to make it ready to support up to 80 channels. We are going to change optical parameters and insert new elements in nodes databases.

During the work, no traffic impact is expected as circuits will be switched to protection path.
Date:  Wednesday, 22 May 2019, 17:16PM
Reason for closing:  Done
Comment by OVH - Wednesday, 22 May 2019, 05:28AM

We are currently working on this line but are facing an issue on the Globalswitch node. Because of a defective amplifier card, we are not able to bring back up circuits on the line Paris GSW <> SBG. All circuits running on this route are currently on protection path with no impact.
A spare card and a field engineer are being sent on site in order to replace the card and should be there in a couple of hours.


Comment by OVH - Wednesday, 22 May 2019, 11:06AM

Field is currently on site. Card is being replaced.


Comment by OVH - Wednesday, 22 May 2019, 17:16PM

Card has been replaced but this did not fix the issue. TAC has kept investigating and finally isolated an issue on the Bellot amplifier site where the new optical parameters have not been properly applied and were preventing the APC (Automatic Power Control) to run on the line and preventing circuits to come back up.
After some manipulatons, we succeded in applying those parameters, run APC and bring all the 26 circuits back online.

Redundancy is now re-estesblished but there are still some things to clarify about why those parameters have not been applied. We are also experiencing a default on one of the amplifier in Bellot that we'll need to replace later. It is not impacting and not urgent, we'll organise that in a few days.

Work is now completed 17h+ after we started yesterday at 23:30.